OpenStudy (anonymous):

i need help! i need to know how to solve this. sqrt(7^5u^12v^20w^65x^80

OpenStudy (anonymous):

\[\sqrt{7^{5}u ^{12} v ^{20} w ^{65} z ^{80}}\]

OpenStudy (anonymous):

divide each exponent by the index, which is 2 so for example 2 goes in to 5 twice, with a remainder of 1 therefore \[\sqrt{7^5}=7^2\sqrt{7}\]

OpenStudy (anonymous):

2 goes in to 12 six times with no remainder so \[\sqrt{u^{12}}=u^6\]

OpenStudy (anonymous):

similarly 2 goes in to 20 ten times with no remainder, so \[\sqrt{v^{20}}=v^{10}\]

OpenStudy (anonymous):

and so on

OpenStudy (anonymous):

so everythin has to be divided into 2?

OpenStudy (anonymous):

yes divide all the exponents by 2 if there is a remainder of 1, i.e. if the exponent is odd, one stays in and the rest comes out of radical

OpenStudy (anonymous):

ok so \[\sqrt{w ^{65}}= w \sqrt{w ^{32}}\]

OpenStudy (anonymous):

no not quite

OpenStudy (anonymous):

how many times does two go in to 65?

OpenStudy (anonymous):

oh i see the mistake you have it backwards \[\sqrt{w^{65}}=w^{32}\sqrt{w}\]

OpenStudy (anonymous):

ohhhh ok. so i have to flip it. but in a way i was correct?

OpenStudy (anonymous):

\(w^{32}\) comes out of the radical, one \(w\) stays in

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

x^40\[\sqrt{x}\]

OpenStudy (anonymous):

\[x^{40}\sqrt{x}\]

OpenStudy (anonymous):

how many times does 2 go in to 40?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

is there a remainder?

OpenStudy (anonymous):

im dividing it 40 by 2

OpenStudy (anonymous):

yes divide 40 by 2 you get 20 but no remainder right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

this means \[\sqrt{x^{40}}=x^{20}\]

OpenStudy (anonymous):

since there is no remainder, nothing is left inside the radical it all comes out

OpenStudy (anonymous):

\[^{20}\sqrt{x}\]

OpenStudy (anonymous):

satellite, your awesome. :)

OpenStudy (anonymous):

x

OpenStudy (anonymous):

(blush)

OpenStudy (anonymous):

@mperez1312 if there is no remainder, nothing stays inside the radical

OpenStudy (anonymous):

for example \(\sqrt{2^6}=2^3\) nothing is left in the radical we see that \(\sqrt{2^6}=\sqrt{64}=8\) and \(2^3=8\) there is no radical because 2 goes in to 6 evenly

OpenStudy (anonymous):

ok perfect! i got it!

OpenStudy (anonymous):

thank you for your help!